# CODI: Compressing Chain-of-Thought into Continuous Space via Self-Distillation

### v1

@@ Appendix A Implementation Details
For iCoT training of GPT-2, we use a learning rate of 5e-5 and train for 100 epochs, removing 4 tokens per epoch for GSM8k-Aug-NL. For iCoT training of LLaMA-1b, we use a learning rate of 1e-5 and train for 50 epochs, removing 8 tokens per epoch for GSM8k-Aug and 16 tokens per epoch for GSM8k-Aug-NL.
@@ Appendix C Datasets
C.1 Examples
GSM8k-Aug* Question = "Jen shared a pack of chocolates among her friends. She gave 20% to Lucy, 30% to Sarah and the remaining were shared equally among four others. If the pack contained 100 chocolates, how many chocolates were each of the four others getting?" CoT = "The total percentage given to Lucy and Sarah is 20% + 30% = 50%. So, the remaining percentage that was shared among the others is 100% - 50% = 50%. The total number of chocolates shared among the others is 100 * 50 / 100 = 50 chocolates. So, each of the four others received 50 / 4 = 12.5 chocolates." Answer = "12.5"
@@ C.1 Examples
C.2 Statistics
Training Dataset Num. Data Avg. CoT Tokens
GSM8k-Aug 385,620 20.3
GSM8k-Aug* 384,625 49.0
Table A1: Training data statistics.
Evaluation Benchmark Num. Data
GSM8k 1319
SVAMP 1000
GSM-Hard 1319
MultiArith 500
Table A2: Evaluation Benchmark statistics.
Appendix D Performance on Training Data
The training accuracy of CODI and CoT-SFT is shown in Table A3.
GPT-2 LLaMA1b
CoT-SFT 99.1% 98.9%
CODI (Ours) 81.5% 95.0%
Table A3: Training Accuracies on GSM8k-Aug.

### v2

@@ Appendix A Implementation Details
For iCoT training of GPT-2, we use a learning rate of 5e-5 and train for 100 epochs, removing 4 tokens per epoch for GSM8k-Aug-NL. For iCoT training of LLaMA-1b, we use a learning rate of 1e-5 and train for 50 epochs, removing 8 tokens per epoch for GSM8k-Aug and 16 tokens per epoch for GSM8k-Aug-NL. LoRA is not used during training.
For Coconut training of GPT-2, we use a learning rate of 1e-4 and train for 25 epochs without continuous tokens and 25 epochs with continuous tokens (50 epochs in total). For iCoT training of LLaMA-1b, we use a learning rate of 1e-5 and train 5 epochs for both stages (10 epochs in total). LoRA is not used during training.
@@ Appendix C Datasets
C.1 Statistics
Training Dataset Num. Data Avg. CoT Tokens
GSM8k-Aug 385,620 20.3
GSM8k-Aug-NL 384,625 49.0
CommonsenseQA-CoT 8,096 85.0
Table A1: Training data statistics.
Evaluation Benchmark Data Size
GSM8k 1,319
SVAMP 1,000
GSM-Hard 1,319
MultiArith 500
CommonsenseQA 1,221
Table A2: Evaluation Benchmark statistics.
C.2 Examples
@@
GSM8k-Aug-NL Question = "Jen shared a pack of chocolates among her friends. She gave 20% to Lucy, 30% to Sarah and the remaining were shared equally among four others. If the pack contained 100 chocolates, how many chocolates were each of the four others getting?" CoT = "The total percentage given to Lucy and Sarah is 20% + 30% = 50%. So, the remaining percentage that was shared among the others is 100% - 50% = 50%. The total number of chocolates shared among the others is 100 * 50 / 100 = 50 chocolates. So, each of the four others received 50 / 4 = 12.5 chocolates." Answer = "12.5"
CommonsenseQA-CoT Question: "The sanctions against the school were a punishing blow, and they seemed to what the efforts the school had made to change? Choices: A: ignore B: enforce C: authoritarian D: yell at E: avoid" CoT = "The context of the sentence indicates that the sanctions are undermining or dismissing the efforts made by the school to change. The word "ignore" fits best here, as it conveys the sense of the sanctions not acknowledging the school’s efforts." Answer = "A"
@@
Appendix D CODI’s Pattern Learning
GPT-2 No-CoT-SFT CODI Coconut Res Op-Res
Accuracy 19.1% 43.7% 34.1% 34.0% 35.7%
Table A3: Comparison of GPT-2 finetuned on two datasets derived from CODI’s decoded thoughts. Res: using intermediate results as CoT. Op-Res: using intermediate operators and results as CoT.
Given that CODI’s continuous thoughts can often be decoded into intermediate results, it raises a question: is CODI effectively equivalent to a GPT-2 fine-tuned on a dataset containing CODI’s decoded patterns? We created a dataset containing only intermediate results (e.g., “CoT: 20, 7, 27. Result: 9” translated from the case study in Figure 6). Additionally, since some cases of CODI show decoded operators like ‘$\times$’ and ‘--’ interleaved with intermediate results, we also create a synthetic CoT dataset that includes both operators and results (e.g., “CoT: $\times$, 20, $\times$, 7, $+$, 27. Result: 9”). As shown in Table A3, while models trained on the two synthetic datasets outperform the No-CoT-SFT baseline, they perform much worse compared to CODI, though perform on par with Coconut. These result suggest that CODI learns richer information from the teacher task through distillation than pure imitation on language-level intermediate results alone, highlighting the advantages of our training framework.